OpenStudy (anonymous):

\[\sqrt[9]{x}^{7} \over \sqrt[9]{x}^{7}\] \[\sqrt[9]{x}^{7} \over \sqrt[9]{x}^{7}\] @Mathematics

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

x^(7/9) /x^(7/9) = x^(7/9 - 7/9) = x^0 = 1

OpenStudy (anonymous):

the second ^7 is suppose to be a ^5

OpenStudy (anonymous):

^(7/9) /x^(5/9) =x^(7/9 - 5/9) = x^(2/9)

OpenStudy (anonymous):

\[\sqrt[9]{x ^{2}}\]